# Random pairing MLE for estimation of item parameters in Rasch model

### v1

@@ Appendix F Auxiliary lemmas
Lemma 11 (A quantitative version of Sylvester’s law of inertia, [Ost59]).
@@
If $w_{ij}>0$ for all $(i,j)\in\mathcal{G}$ and $\mathcal{G}$ is a connected graph, then $\bm{L}_{w}$ is rank $m-1$, $\bm{L}_{w}\bm{1}_{m}=\bm{0}_{m}$ and $\bm{L}_{w}^{\dagger}\bm{1}_{m}=0$. Moreover for any $i\in[n-1]$, $\lambda_{i}(\bm{L}_{w}^{\dagger})=\lambda_{n-i}(\bm{L}_{w})$.
@@
The fact that $\bm{L}_{w}$ is rank $m-1$ when $\mathcal{G}$ is connected is well-known. See e.g. [Spi07] for reference. Since $\bm{L}_{w}$ is a real symmetric matrix, it has an eigendecomposition $\bm{L}_{w}=\bm{U}\bm{\Sigma}\bm{U}^{\top}$ and then $\bm{L}_{w}^{\dagger}=\bm{U}\bm{\Sigma}^{\dagger}\bm{U}^{\top}$. The rest follows from this decomposition and the form of $\bm{L}_{w}$.∎

### v2

@@ Appendix E Auxiliary lemmas
Lemma 13 (A quantitative version of Sylvester’s law of inertia, [35]).
@@
If $w_{ij}>0$ for all $(i,j)\in\mathcal{G}$ and $\mathcal{G}$ is a connected graph, then $\bm{L}_{w}$ is rank m1m-1, $\bm{L}_{w}\bm{1}_{m}=\bm{0}_{m}$ and $\bm{L}_{w}^{\dagger}\bm{1}_{m}=0$. Moreover for any $i\in[n-1]$, $\lambda_{i}(\bm{L}_{w}^{\dagger})=\lambda_{n-i}(\bm{L}_{w})$.
@@
The fact that $\bm{L}_{w}$ is rank m1m-1 when $\mathcal{G}$ is connected is well-known. See e.g. [41] for reference. Since $\bm{L}_{w}$ is a real symmetric matrix, it has an eigendecomposition $\bm{L}_{w}=\bm{U}\bm{\Sigma}\bm{U}^{\top}$ and then $\bm{L}_{w}^{\dagger}=\bm{U}\bm{\Sigma}^{\dagger}\bm{U}^{\top}$. The rest follows from this decomposition and the form of $\bm{L}_{w}$.∎
E.1 Proof of Fact 1
Expanding the probability of the random events, we have
[…]
$=\frac{e^{\zeta_{t}^{\star}}e^{\theta_{j}^{\star}}}{(e^{\zeta_{t}^{\star}}+e^{\theta_{i}^{\star}})(e^{\zeta_{t}^{\star}}+e^{\theta_{j}^{\star}})}\cdot$
$\qquad\left(\frac{e^{\zeta_{t}^{\star}}e^{\theta_{j}^{\star}}}{(e^{\zeta_{t}^{\star}}+e^{\theta_{i}^{\star}})(e^{\zeta_{t}^{\star}}+e^{\theta_{j}^{\star}})}+\frac{e^{\theta_{i}^{\star}}e^{\zeta_{t}^{\star}}}{(e^{\zeta_{t}^{\star}}+e^{\theta_{i}^{\star}})(e^{\zeta_{t}^{\star}}+e^{\theta_{j}^{\star}})}\right)^{-1}$
$=\frac{e^{\zeta_{t}^{\star}}e^{\theta_{j}^{\star}}}{e^{\zeta_{t}^{\star}}(e^{\theta_{i}^{\star}}+e^{\theta_{j}^{\star}})}=\frac{e^{\theta_{j}^{\star}}}{e^{\theta_{i}^{\star}}+e^{\theta_{j}^{\star}}}.$
Now consider $\mathbb{P}[X_{ti}\neq X_{tj}]$, we have
Let $f:[1/\kappa_{2},\kappa_{2}]^{2}\rightarrow\mathbb{R}$ defined by
$f(a,b)\coloneqq\frac{a+b}{(1+a)(1+b)}.$
Its partial derivatives are